OpenStudy (anonymous):

Simplify the expression (8/125)^-2/3

jimthompson5910 (jim_thompson5910):

\[\left(\frac{8}{125}\right)^{-\frac{2}{3}}\] \[\left(\frac{125}{8}\right)^{\frac{2}{3}}\] \[\left(\sqrt[3]{\frac{125}{8}}\right)^2\] \[\left(\frac{5}{2}\right)^2\] \[\frac{25}{4}\] So \[\left(\frac{8}{125}\right)^{-\frac{2}{3}}=\frac{25}{4}\]